OpenStudy (anonymous):

A 20.0 gram bullet is shot horizontally into a 2.00 kg block and quickly becomes imbedded in the block. After the impact, the block slides along a rough horizontal surface until it comes to a stop after moving 85 cm. The coefficient of kinetic friction between the block and the surface is μk = 0.45. How much energy is dissipated in the block during the impact?

OpenStudy (anonymous):

Stopping distance = d = V^2 / 2 * u * g d = distance = 0.85m V = velocity of object before friction applied u = kinetic friction coeff = 0.45 g = gravity = 9.81 m/s^2 d = V^2 / (2 * u * g) 0.85m = V^2 / 2 * 0.45 * 9.81 m/s^2 V^2 = (2 * 0.45 * 9.81 m/s^2) * 0.85m V = sq rt [(2 * 0.45 * 9.81 m/s^2) * 0.85m] Velocity of block prior to deceleration by friction: V = 2.73 m/s Now take kinetic energy knowing the block had to be going 2.73 m/s before it slid to stop: KE = 1/2 mv^2 KE = 1/2 (0.02kg + 2kg) (2.73 m/s)^2 KE = 7.5 Joules of energy lost in impact

OpenStudy (anonymous):

Thank you so much! I totally did not understand what they were asking.